a bucket is 20cm in diameter at the open end, 12cm in diameter at the bottom and 16cm deep. to what depth would the bucket fill a cylindricial tin 28cm in diameter
1
Expert's answer
2010-12-20T16:39:11-0500
We have to evaluate the volume of the bucket:
V = 1/3 π H (R2 + Rr +r2) = 1/3 π *16 (202 +20*12 +122) = 13136,0461 cm2
The volume of the cylindricial tin is

V = π R2 h = π* 282*h = 13136,0461;
h = 5.33 cm


Answer: the depth is 5.33 cm.